OpenStudy (pottersheep):

If u = x^1/4 , why would u^2 = 1/2?

OpenStudy (jamesj):

This is what you're asking across this post and the previous one: \[u = x^{1/4} \implies u^2 = (x^{1/4})^2 = x^{2/4} = x^{1/2}\]